A ball is launched directly upward from ground lev
Answer #1

(a)

Use kinematic equation,

v = u + at

where v is final speed ,u is initial speed ,a = -g and t = time

0 = 22 m/s - 9..8 t

Solve for t

t = 2.24 s

This is time of ascent. Total time in air is twice the time ascent.

2t = 4.48 s

Round off into 2 significant digits

2 t = 4.5 s

------------------------------------------------------------------------------------------------------------------------------------

(b)

From kinematic equation

v^2 = u^2 +2aS

0 = ( 22 m/s ) ( 22 m/s ) -2 (9.8 ) S

Solve for S

S = 24.44 m

Round off into 2 significant digits

  S = 24 m

----------------------------------------------------------------------------------------------------------------------------------------

(c)

From kinematic equation

S = ut +1/2at2

24.44 m = 22 t - 1/2 *9.8 *t2

4.9 t2 -22 t +24.44 =0

Solve quadratic equation for t

t = 2.01 s ( first time ) and 2.47 ( second time )

Round off into 2 significant digits

t = 2.0 s ( first time ) and 2.5 s ( second time )